OpenStudy (anonymous):

Differentiate the following by the first principle : f(x) = 2x + 3

OpenStudy (callisto):

It's not differential equation... ------------------------------------------ \[f(x) = 2x + 3\]\[f'(x) = lim_{h\rightarrow 0}\frac{2(x+h) +3 - (2x+3)}{h}\] Can you simplify 2(x+h) +3 - (2x+3)?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

What do you get?

OpenStudy (anonymous):

the ans is 2

OpenStudy (callisto):

I know the answer is 2. I was asking what have you got after simplifying 2(x+h) +3 - (2x+3).

OpenStudy (anonymous):

ans iz 2

OpenStudy (anonymous):

f(x)=2x + 2h +3 - 2x-3/h =2h/h =2 \[\prime(x)\]

OpenStudy (callisto):

Almost correct. It's just the problem of notation. f(x) = 2x+3 \[f'(x)\]\[ = lim_{h\rightarrow 0}\frac{2(x+h) +3 - (2x+3)}{h}\]\[= lim_{h\rightarrow 0}\frac{2x+2h +3 - (2x+3)}{h}\]\[=lim_{h\rightarrow 0}\frac{2h}{h}\]\[=2\]